Solve 2x^-2-x^-1-1=0
Answer by RAY100(1637)

2 x^-2 -x^-1 -1 =0
.
2/x^2 -1/x^1 -1 =0
.
multiply thru by x^2
.
2 -x -x^2 =0
.
mult thru by -1
.
x^2 +x -2 =0
.
(x+2)(x-1) =0
.
x= 1, -2
.
check
(1) 2*1^-2 - (1)^-1 -1 =0
2 -1-1=0,,,,,ok
.
(-2),,,,2*(-2)^-2 -(-2)^-1 -1 =0
.5 +.5 -1 =0,,,,,,,,,ok